Can someone help me solve this?
Answer the following question.
Solve the system by the substitution method.
x-y=6
y=-2-x

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

y=-2-x
x-y=6
x-(-2-x)=6
x+2+x=6
2x=4
x=2
------------
2-y=6
-y=4
y=-4
-----------
-4=-2-2
-4=-4

Answer by stanbon(75887)

Solve the system by the substitution method.
x-y=6
y=-2-x
--------
Substitute for "y" and solve for "x"::
x - (-2-x) = 6
x +2 + x = 6
2x = 4
x = 2
--------------
Solve for "y":
y = -2-x
y = -2-2
y = -4
----------------
Cheers,
Stan H.
----------------